# Collatz-Weyl Generators: High Quality and High Throughput Parameterized Pseudorandom Number Generators

### v2

@@ 1 INTRODUCTION
It is well known that no arithmetical method is available to automatically generate random numbers. The pseudorandom number generator (PRNG) is a deterministic algorithm that generates numbers that approximate to the properties of random numbers. Designing a good pseudorandom number generator is always a bit of an art, not strict mathematics. As von Neumann said, “we are dealing with mere ‘cooking recipes’ for making digits; they can probably not be justified, but should merely be judged by their results” von1951various .
@@
In this paper, the author proposes a scheme to create high-quality PRNGs. The proposed scheme is based on functions with strong mixing properties, called the generalized Collatz functions (GCFs). To eliminate the problem of short cycles we cmobine them with Weyl sequences. We focus on simplified GCFs to maximize efficiency in implementations of this scheme. However, we think that more complex GCFs are also suitable for the generation of pseudorandom numbers, especially in cryptography, owing to a wide parameter space and the non-invertibility [^1] Note that when we use the term “non-invertibility,” we refer only to graph structures of the state space of the generator. For instance, there may be several paths to reverse the state of such a generator. We do not refer to non-invertibility in the cryptographic sense at any point in this paper. However, we suggest that the non-cryptographic non-invertibility of some of the presented schemes may remain consistent with non-invertibility in the cryptographic sense. We leave this issue as a hypothesis, and do not analyze it in this paper. of such mappings as this may result in a resistance to backtracking in specific schemes fischer2012backtracking . Furthermore, past research has shown that such functions can be associated with factorization as well as the discrete logarithm problem, where this makes them suitable for cryptographic applications thakkar2016additive . We leave this issue open for future research.
@@ 2.1 Generalized Collatz Sequences
Generalized Collatz sequences have been considered by several researchers. Conway conway1972unpredictable and, subsequently, Kurtz and Simon kurtz2007undecidability considered the following generalizations (generalized Collatz functions):
@@ 2.2 General Scheme
Definition 2 allows for the creation of a wide class of State Functions that can be implemented in PRNGs. However, because such functions usually define non-invertible pseudorandom mappings, we mix their states with Weyl sequences–counters modulo $2^{w}$, where each increment is an odd number marsaglia2003xorshift ; widynski2017middle –to provide the generator with a guaranteed minimum period. An additional step involving mixing coefficients and the outputs is also usually required to yield results of better quality.
@@ 2.3 Algorithms
$C(x)=\left\{\begin{array}[]{cl}\frac{3x+1}{2}&\mbox{if}~{}x\equiv 1~{}~{}(\bmod~{}2),\\ \\ \frac{x}{2}&\mbox{if}~{}~{}x\equiv 0~{}~{}(\bmod~{}2).\end{array}\right.$
@@
$C(x)=\left\{\begin{array}[]{cl}x=x+((x+1)\gg 1)&\mbox{if}~{}x\equiv 1~{}~{}(\bmod~{}2),\\ \\ x=x\gg 1&\mbox{if}~{}~{}x\equiv 0~{}~{}(\bmod~{}2).\end{array}\right.$
@@ 3 PERIOD
Collatz-Weyl Generators can be considered to be a combination of two sequences: one defined by a non-invertible pseudorandom mapping, and the other being a Weyl sequence of an irreducible, minimum-period length. The cycle length of a typical combined generator is the least common multiple of the cycle lengths of the individual generators fog2015pseudo . This technique for extending the period of the generator has been known since Marsaglia marsaglia2003xorshift , and was used by Widynski widynski2017middle . However, the Marsaglia design involves two separate generators that are combined in the output. In our design here, the state of the generator, and not merely the output, is combined with the Weyl sequence in a similar manner to that in Widynski’s middle-square Weyl generator. Therefore, there is no least common multiple of two separate generators because we are dealing with only one generator with a dual structure. However, we may estimate period of the non-invertible pseudorandom mapping and the irreducible Weyl sequence. Because the period of an n-bit Weyl sequence is well known, the minimum period is guaranteed to be $2^{n}$. The period of the non-invertible pseudorandom sequence may be approximated from the theory of random mappings flajolet1989random . The expected path to reach the cycle and the expected cycle length are given by the same formula,
@@
, for an n-bit state generator. On the whole, we start to see repeated patterns after
@@
iterations. If we had a pseudorandom number generator composed of three 64-bit substates based on a non-invertible pseudorandom sequence, like CWG64, then we would calculate the number of iterations after which we expect repetitions by substituting $n=196$. However, the non-invertible state of the CWG64 generator is composed of 128 bits (64-bit substate $x\gg 1$ and 64-bit substate $(a+=x)|1$ are multiplied), and the 64-bit Weyl substate has irreducible cycle lenght equal to $2^{64}$. Therefore the expected cycle length and the expected path to reach the cycle will be a multiple of the Weyl sequence
$\sqrt{\frac{\pi}{8}\cdot 2^{128}}\cdot 2^{64}$
. We thus start to see repeats after
$2\cdot\sqrt{\frac{\pi}{8}\cdot 2^{128}}\cdot 2^{64}\sim 2^{128.33}$
numbers in the CWG64 generator. By analogy, we start to see repeats in the CWG128-64 generator and the CWG128 generator after
$2\cdot\sqrt{\frac{\pi}{8}\cdot 2^{128}}\cdot 2^{64}\sim 2^{128.33}$
@@
$2\cdot\sqrt{\frac{\pi}{8}\cdot 2^{256}}\cdot 2^{128}\sim 2^{256.33}$
@@
Because the $2^{n}$ cycle length of the n-bit Weyl sequence is guaranteed, a cycle shorter than the length of the Weyl sequence cannot exist in our generator because if so, we would be able to initialize the generator so that it would start in such a cycle from the first iteration such that the minimum guaranteed period would not be met. However, arbitrarily short paths to reach the cycles may occur. In particular, we can imagine an initialization that starts the generator in a certain cycle, in which case the length of the path to reach the cycle would be zero.
In summary, the minimum guaranteed period of CWG64 under every uniquely initialized Weyl sequence (there are $2^{63}$ such initializations) is $2^{64}$, and we expect to see repeats after $\sim 2^{128.33}$ generated numbers. That of CWG128-64 under every uniquely initialized Weyl sequence (there are $2^{63}$ such initializations) is $2^{64}$, and we expect to see repeats after $\sim 2^{128.33}$ generated numbers. The minimum guaranteed period of CWG128 under every uniquely initialized Weyl sequence (there are $2^{127}$ such initializations) is $2^{128}$, and we expect to see repeats after $\sim 2^{256.33}$ generated numbers.
@@ 4 UNIFORMITY
Uniformity is a property that implies that a PRNG returns each value in the range of possible results with equal probability. The Weyl sequence provides not only a minimum guaranteed period, but also uniformity. Widynski gave simple proof of this widynski2017middle .
@@
Let $z=(x>>1)*((a+=x)|1)$ be a random but not necessarily uniform bit stream, and let $weyl$ be a uniform but not necessarily random bit stream. Then, $z\oplus weyl$ is a uniform bit stream.
@@
Given an arbitrary element $y$ in $[0,2^{n}-1]$, we compute the probability of $y$, where $y=z\oplus weyl$. The probability distribution of the sum of two random variables is the convolution of their individual distributions. This also holds for modulo-2 addition. Therefore:
$P(y)=\sum_{z=0}^{2^{n}-1}(P(z)\cdot P(weyl))$
@@
There are $2^{n}$ possible outputs in the n-state Collatz-Weyl Generator, each of which has a probability of occurrence of $\frac{1}{2^{n}}$. Such a generator produces uniformly distributed values of $x=z\oplus weyl$, and $a$ is thus also uniform as it is simply the modulo sum of uniform numbers. The output of Collatz-Weyl Generators is therefore a bitwise shift and an XOR operation on two uniform variables, and thus is also uniform. Note that this proof refers only to the occurrence of each value in the range of possible results with an equal probability. This does not mean that each value in the range of $2^{n}$ outputs occur exactly once. Furthermore, even if our generator were based on a perfectly random bit stream, we expect a particular number of repetitions in the values of the generated numbers due to the birthday problem wiki:birthday .
@@ 5 BIRTHDAY PROBLEM
As we show below, Collatz-Weyl Generators pass statistical tests above the guaranteed minimum period. However, because they are based on pseudorandom mappings, there is no risk of flaws due to the birthday problem such that they can be used in the range of the full period, rather than its square root as in the case of typical PRNGs caflisch1998monte ; larcher1998random ; gentle2003random ; o2015pcg . The uniformity of the generators proven in the previous section does not guarantee that all outputs will have occurred exactly the same number of times once a generator has completed a full period, but implies only that a generator returns each value in the range of possible results with an equal probability. However, some numbers may appear twice or more by chance, as in case of the generation of true random numbers. Even if we initialized a generator so that it did not go through a path to reach the cycle but started at a particular cycle (in the worst case, such a cycle may have only the length of the underlying Weyl sequence), numbers in this cycle usually do not form a one-cycle permutation as in the LCG generator, for example, in which each number occurs exactly once in a period (which leads to a deviation from the true randomness detectable by the birthday test). We expect cycles composed of random numbers in CWGs that are chosen uniformly from the range of possible outputs, and therefore some numbers may repeat. Of course, we need to remember that each pseudorandom number generator has an internal bias, and the proposed generators simulate only randomness. However, PRNGs based on random mappings seem to be the closest we can approach to emulating true randomness. This holds as long as they do not contain systematic bias, and provide guarantees on the fundamental characteristics of the generator related to the confidence that we can have in it, such as its period and uniformity.
@@ 6.1 Initialization
The increment $s$ must be odd to ensure the maximum period of the Weyl sequence. All the other variables can be initialized by using any number. We suggest seeding $s$ by an odd number and setting all other elements to zero for the sake of simplicity, except CWG128-64 where $x$ should also be non-zero.
To ensure the complete independence of the $2^{64}$ streams, the first 48 and 96 states of each stream should be skipped in the CWG64 and CWG128-64, and the CWG128 generators, respectively. Otherwise, minor correlations may occur in the first outputs of the successively initialized streams, especially if they are initialized by using the numbers $1,3,5,...$ or another counter. This is because CWG generators initialized with similar but not the same states need a certain number of iterations to decorelate. For this purpose, the user may call the initializer (Figures 5, 6, and 7). Every stream can be initialized with any 64-bit odd number by using the initializer().
@@
For faster initialization, the user may apply the Splitmix64 steele2014fast and Splitmix63 generators (a modified Splitmix64 generator), and use their outputs to serially fill the values of $x$ and $s$. The implementations of the Splitmix64 and Splitmix63 generators in C++ are shown in Figures 8 and 9. We recommend using the method shown in Figure 10 to serially seed the state of CWG64, this makes it possible to initialize $2^{63}$ streams, the method presented in Figure 11 to serially seed the state of CWG128-64, this makes it possible to initialize $2^{63}$ streams, and to serially seed the state of CWG128 using the method shown in Figure 12, this makes it possible to initialize $2^{63}$ streams.
@@
The Splitmix63 generator is required to obtain a unique and non-repetitive sequence of 63-bit numbers that are needed to initialize odd Weyl increments. This would not be possible with the Splitmix64 generator [^2] There is a low probability of obtaining identical numbers, i.e., a collision, whereby 64-bit numbers are converted into 63-bit numbers.. Research has shown matsumoto2007common that initialization performed by using a generator that is radically different from the one initialized enables the avoidance of a correlation on similar seeds. A similar method has been used in the Xoroshiro family of PRNGs blackman2021scrambled . We think that this will also work for our generators.
@@ 7 STATISTICAL TESTS
All the proposed generators passed the NIST test suite, Dieharder, TestU01 (Big Crush), and PractRand up to 8 TB rukhin2001statistical ; brown2018dieharder ; l2007testu01 ; doty10practically . Since standard versions of CWG generators may have to big states to fail and so applied tests would not be able to identify their limitations too-big-to-fail ; pcg-passes-practrand ; o2015pcg , we examined scaled down versions of CWG64 (CWG32 and CWG16) and CWG128-64 (CWG64-32 and CWG32-16). All scaled-down versions passed PractRand above the limit of their minimum guaranteed periods, which confirms the quality of their full-scale versions.
@@ 7.1 Methodology
We executed 200 complete runs of the three proposed generators to verify the dependence of random numbers from the same sequence (intra-stream correlation), and 400 complete runs to verify their inter-stream correlations salmon2011parallel ; tan2021thundering .
We used two methods to verify the dependence between the separate sequences. The first was adapted from work by Li et al. li2013software . We first interleaved 4, 8, 128, and 1024 sequences into a single sequence and then evaluated the resulting sequence. 100 complete runs were performed for initialization by the initalizator() and 100 complete runs for initialization by the Splimixes. In the second method we tested consecutive firsts, seconds, fourths, and 10ths of the outputs (so we created one stream composed of the first outputs of serially initialized generators, then tested a stream composed of the second outputs of serially initialized generators, and so on). 100 complete runs for initialization by the initializer() and 100 complete runs for initialization by the Splimixes were performed.

### v4

@@ 1 INTRODUCTION
It is well known that no arithmetical method is available to automatically generate random numbers. The pseudorandom number generator (PRNG) is a deterministic algorithm that generates numbers that approximate to the properties of random numbers. Designing a good pseudorandom number generator is always a bit of an art, not strict mathematics. As von Neumann said, “we are dealing with mere ‘cooking recipes’ for making digits; they can probably not be justified, but should merely be judged by their results” [1].
@@
In this paper, the author proposes a scheme to create high-quality PRNGs. The proposed scheme is based on functions with strong mixing properties, called the generalized Collatz functions (GCFs). To eliminate the problem of short cycles we cmobine them with Weyl sequences. We focus on simplified GCFs to maximize efficiency in implementations of this scheme. However, we think that more complex GCFs are also suitable for the generation of pseudorandom numbers, especially in cryptography, owing to a wide parameter space and the non-invertibility [^1] Note that when we use the term “non-invertibility,” we refer only to graph structures of the state space of the generator. For instance, there may be several paths to reverse the state of such a generator. We do not refer to non-invertibility in the cryptographic sense at any point in this paper. However, we suggest that the non-cryptographic non-invertibility of some of the presented schemes may remain consistent with non-invertibility in the cryptographic sense. We leave this issue as a hypothesis, and do not analyze it in this paper. of such mappings as this may result in a resistance to backtracking in specific schemes [2]. Furthermore, past research has shown that such functions can be associated with factorization as well as the discrete logarithm problem, where this makes them suitable for cryptographic applications [3]. We leave this issue open for future research.
@@ 2.1 Generalized Collatz Sequences
Generalized Collatz sequences have been considered by several researchers. Conway [4] and, subsequently, Kurtz and Simon [5] considered the following generalizations (generalized Collatz functions):
@@ 2.2 General Scheme
Definition 2 allows for the creation of a wide class of State Functions that can be implemented in PRNGs. However, because such functions usually define non-invertible pseudorandom mappings, we mix their states with Weyl sequences–counters modulo $2^{w}$, where each increment is an odd number [6, 7]–to provide the generator with a guaranteed minimum period. An additional step involving mixing coefficients and the outputs is also usually required to yield results of better quality.
@@ 2.3 Algorithms
$C(x)=\left\{\begin{array}[]{cl}\frac{3x+1}{2}&\mbox{if}~{}x\equiv 1~{}~{}(\bmod~{}2),\\ {}{}{}\hfil\\ \frac{x}{2}&\mbox{if}~{}~{}x\equiv 0~{}~{}(\bmod~{}2).\end{array}\right.$
@@
$C(x)=\left\{\begin{array}[]{cl}x=x+((x+1)\gg 1)&\mbox{if}~{}x\equiv 1~{}~{}(\bmod~{}2),\\ {}{}{}\hfil\\ x=x\gg 1&\mbox{if}~{}~{}x\equiv 0~{}~{}(\bmod~{}2).\end{array}\right.$
@@ 3 PERIOD
Collatz-Weyl Generators can be considered to be a combination of two sequences: one defined by a non-invertible pseudorandom mapping, and the other being a Weyl sequence of an irreducible, minimum-period length. The cycle length of a typical combined generator is the least common multiple of the cycle lengths of the individual generators [8]. This technique for extending the period of the generator has been known since Marsaglia [6], and was used by Widynski [7]. However, the Marsaglia design involves two separate generators that are combined in the output. In our design here, the state of the generator, and not merely the output, is combined with the Weyl sequence in a similar manner to that in Widynski’s middle-square Weyl generator. Therefore, there is no least common multiple of two separate generators because we are dealing with only one generator with a dual structure. However, the period of the Weyl sequence is well known, and we may estimate period of the non-invertible pseudorandom mapping, to determine the overall period of the generator. Because the period of an n-bit Weyl sequence is irreducible, the minimum period of every CWG generator is guaranteed to be $2^{n}$. The period of the non-invertible pseudorandom sequence may be approximated from the theory of random mappings [9]. The expected path to reach the cycle and the expected cycle length are given by the same formula
@@
for an n-bit state generator. On the whole, we start to see repeated patterns after
@@
iterations. If we had a pseudorandom number generator composed of two 64-bit substates based on a non-invertible pseudorandom sequence, then we would calculate the number of iterations after which we expect repetitions by substituting $n=128$. But here we must also consider the irreducible substate $weyl$. In CWG64 generator, only the non-invertible part of the state $x$ is xored with 64-bit $weyl$ substate, and $a$ is not. Hence, only 64 bits of the generator state will preserve the non-invertible pseudorandom mapping properties, while the other 64 bits would follow the period of the Weyl sequence. Therefore, the generator period will be a product of the Weyl sequence period and the expected value of path and cycle length of the 64-bit non-invertible random mapping:
$\sqrt{\frac{\pi}{8}\cdot 2^{64}}\cdot 2^{64}$
. Hovewer, this does not imply that $x$ will have exactly the $weyl$ period, nor that $a$ will follow the period predicted by pseudorandom mapping statistics, as they are both mixed together. Their periods will be given by the same formula and synchronized, especially because substate $a$ is simply the modulo sum of $x$. Consequently it will have a path to the cycle of the same length as $x$ and a cycle perfectly synchronized with $x$. Thus, in overall we expect to see repeats in CWG64 generator after
$2\cdot\sqrt{\frac{\pi}{8}\cdot 2^{64}}\cdot 2^{64}\sim 2^{96.33}$
numbers. By analogy, we start to see repeats in the CWG128-64 generator and the CWG128 generator after
$2\cdot\sqrt{\frac{\pi}{8}\cdot 2^{64}}\cdot 2^{64}\sim 2^{96.33}$
@@
$2\cdot\sqrt{\frac{\pi}{8}\cdot 2^{128}}\cdot 2^{128}\sim 2^{192.33}$
@@
Because the $2^{n}$ cycle length of the n-bit Weyl sequence is guaranteed, a cycle shorter than the length of the Weyl sequence cannot exist in our generator. Every cycle will be a multiple of the Weyl sequence’s cycle length. However, arbitrarily short paths to reach the cycles may occur. In particular, we can imagine an initialization that starts the generator in a certain cycle, in which case the length of the path to reach the cycle would be zero. A generator initialized with the same $s$ or $c[0]$ can have several disjoint cycles. The expected number of components (in our case, this corresponds to the number of cycles, since each component contains one unique cycle) based on the theory of non-invertible random mappings [9] is given by:
$\frac{\ln(2^{n})}{2}$
, where $n$ is equal to $64$ for CWG64 and CWG128-64, and $128$ for CWG128.
In summary, the minimum guaranteed period of CWG64 under every uniquely initialized Weyl sequence (there are $2^{63}$ such initializations) is $2^{64}$, and we expect to see repeats after $\sim 2^{96.33}$ generated numbers. That of CWG128-64 under every uniquely initialized Weyl sequence (there are $2^{63}$ such initializations) is $2^{64}$, and we expect to see repeats after $\sim 2^{96.33}$ generated numbers. The minimum guaranteed period of CWG128 under every uniquely initialized Weyl sequence (there are $2^{127}$ such initializations) is $2^{128}$, and we expect to see repeats after $\sim 2^{192.33}$ generated numbers.
@@ 4 UNIFORMITY
Uniformity is a property that implies that a PRNG returns each value in the range of possible results with equal probability. The Weyl sequence provides not only a minimum guaranteed period, but also uniformity. Widynski gave simple proof of this [7].
@@
Let $x$ be a random but not necessarily uniform bit stream, and let $weyl$ be a uniform but not necessarily random bit stream. Then, $x\oplus weyl$ is a uniform bit stream.
@@
Given an arbitrary element $y$ in $[0,2^{n}-1]$, we compute the probability of $y$, where $y=x\oplus weyl$.
$P(y)=\sum_{x=0}^{2^{n}-1}(P(x)\cdot P(weyl))$
@@
There are $2^{n}$ possible outputs in the n-state Collatz-Weyl Generator, each of which has a probability of occurrence of $\frac{1}{2^{n}}$. Such a generator produces uniformly distributed values of $y=x\oplus weyl$, and $a$ is thus also uniform as it is simply the modulo sum of uniform numbers. The output of Collatz-Weyl Generators is therefore a bitwise shift and an XOR operation on two uniform variables, and thus is also uniform. Note that this proof refers only to the occurrence of each value in the range of possible results with an equal probability. This does not mean that each value in the range of $2^{n}$ outputs occur exactly once. Furthermore, even if our generator were based on a perfectly random bit stream, we expect a particular number of repetitions in the values of the generated numbers due to the birthday problem [10].
@@ 5 BIRTHDAY PROBLEM
As we show below, Collatz-Weyl Generators pass statistical tests above the guaranteed minimum period. However, because they are based on pseudorandom mappings, there is no risk of flaws due to the birthday problem such that they can be used in the range of the full period, rather than its square root as in the case of typical PRNGs [11, 12, 13, 14]. The uniformity of the generators proven in the previous section does not guarantee that all outputs will have occurred exactly the same number of times once a generator has completed a full period, but implies only that a generator returns each value in the range of possible results with an equal probability. However, some numbers may appear twice or more by chance, as in case of the generation of true random numbers. Even if we initialized a generator so that it did not go through a path to reach the cycle but started at a particular cycle (in the worst case, such a cycle may have only the length of the underlying Weyl sequence), numbers in this cycle usually do not form a one-cycle permutation as in the LCG generator, for example, in which each number occurs exactly once in a period (which leads to a deviation from the true randomness detectable by the birthday test). We expect cycles composed of random numbers in CWGs that are chosen uniformly from the range of possible outputs, and therefore some numbers may repeat. Of course, we must remember that each pseudorandom number generator has an internal bias, and the proposed generators simulate only randomness. However, PRNGs based on random mappings seem to be the closest we can approach to emulating true randomness. This holds as long as they do not contain systematic bias, and provide guarantees on the fundamental characteristics of the generator related to the confidence that we can have in it, such as its period and uniformity.
@@ 6.1 Initialization
The increment $s$ must be odd to ensure the maximum period of the Weyl sequence. All the other variables can be initialized by using any number. We suggest seeding $s$ by an odd number and setting all other elements to zero for the sake of simplicity, in single-stream applications.
In multi-stream applications, to ensure the complete independence of the $2^{64}$ streams, the first 48 and 96 states of each stream should be skipped in the CWG64 and CWG128-64, and the CWG128 generators, respectively. Otherwise, minor correlations may occur in the first outputs of the successively initialized streams, especially if they are initialized by using the numbers $1,3,5,...$ or another counter. This is because CWG generators initialized with similar but not the same states need a certain number of iterations to decorelate. For this purpose, the user may call the initializer (Figures 5, 6, and 7). Every stream can be initialized with any 64-bit odd number by using the initializer().
@@
For faster initialization, the user may apply the Splitmix64 [15] and Splitmix63 generators (a modified Splitmix64 generator), and use their outputs to serially fill the values of $x$ and $s$. The implementations of the Splitmix64 and Splitmix63 generators in C++ are shown in Figures 8 and 9. We recommend using the method shown in Figure 10 to serially seed the state of CWG64, this makes it possible to initialize $2^{63}$ streams, the method presented in Figure 11 to serially seed the state of CWG128-64, this makes it possible to initialize $2^{63}$ streams, and to serially seed the state of CWG128 using the method shown in Figure 12, this makes it possible to initialize $2^{63}$ streams.
@@
The Splitmix63 generator is required to obtain a unique and non-repetitive sequence of 63-bit numbers that are needed to initialize odd Weyl increments. This would not be possible with the Splitmix64 generator [^2] There is a low probability of obtaining identical numbers, i.e., a collision, whereby 64-bit numbers are converted into 63-bit numbers.. Research has shown [16] that initialization performed by using a generator that is radically different from the one initialized enables the avoidance of a correlation on similar seeds. A similar method has been used in the Xoroshiro family of PRNGs [17]. We think that this will also work for our generators.
@@ 7 STATISTICAL TESTS
All the proposed generators passed the NIST test suite, Dieharder, TestU01 (Big Crush), and PractRand up to 8 TB [18, 19, 20, 21]. Since standard versions of CWG generators may have to big states to fail and so applied tests would not be able to identify their limitations [22, 23, 14], we examined scaled down versions of CWG64 (CWG32 and CWG16) and CWG128-64 (CWG64-32 and CWG32-16). All scaled-down versions passed PractRand above the limit of their minimum guaranteed periods, which confirms the quality of their full-scale versions.
@@ 7.1 Methodology
We executed 200 complete runs of the three proposed generators to verify the dependence of random numbers from the same sequence (intra-stream correlation), and 400 complete runs to verify their inter-stream correlations [24, 25].
We used two methods to verify the dependence between the separate sequences. The first was adapted from work by Li et al. [26]. We first interleaved 4, 8, 128, and 1024 sequences into a single sequence and then evaluated the resulting sequence. 100 complete runs were performed for initialization by the initalizator() and 100 complete runs for initialization by the Splimixes. In the second method we tested consecutive firsts, seconds, fourths, and 10ths of the outputs (so we created one stream composed of the first outputs of serially initialized generators, then tested a stream composed of the second outputs of serially initialized generators, and so on). 100 complete runs for initialization by the initializer() and 100 complete runs for initialization by the Splimixes were performed.